# How LLMs Learn to Reason: A Complex Network Perspective

### v1

@@ Appendix D CoNet: A Minimal Model for RLVR dynamics
CoNet was originally introduced to model the phenomenon of Learning at Criticality (LaC), which was identified in single-task RLVR training (Cai et al., 2025). The central finding of LaC is that for a model trained on extremely sparse data, generalization performance peaks precisely at a sharp learning transition. The key methodological advance of this paper is to deploy this model in a richer, multi-task setting. This extension is crucial: it transforms CoNet from a model of isolated skill acquisition into an ideal theoretical laboratory for studying the emergent phenomena of skill integration, frustration, and concept web formation. As Fig. 1 shows, this multi-task CoNet stunningly reproduces the key macroscopic signatures of the 1.5B LLM, validating its use for exploring the network-level mechanisms that drive the emergence of reasoning.
@@ Appendix E CoNet Implementation Details
We configured CoNet with a specific set of structural and learning parameters. The underlying concept space was instantiated as a directed random regular graph with $N=800$ nodes, where each node has a uniform out-degree of $k=40$. In this graph, we established a multi-task learning environment consisting of $128$ randomly generated question-answer (Q-A) node pairs. The initial transition parameters $\theta_{ij}$ for all edges were assigned randomly. The use of a multi-task setting with 128 concurrent Q-A pairs is the fundamental mechanism that induces structural competition. While a single task would only require finding one efficient path, the presence of numerous tasks provides competing reinforcement signals. When multiple Q-A paths need to traverse a common node, they may “disagree” on the optimal subsequent step, leading to conflicting policy updates. This microscopic conflict is the origin of the macroscopic “frustration” discussed in the main text, compelling the system to find a shared, globally coherent structure that can arbitrate these conflicts, rather than settling for 128 disconnected “skill islands”.
@@
This advantage-based update functions as a competitive reinforcement mechanism. Only paths performing better than the current average are strengthened, while underperforming paths are implicitly suppressed. This “winner-take-more” dynamic acts as an algorithmic sculpting tool, systematically pruning the vast majority of the initial $k=40$ edges from each node to carve out the sparse, efficient backbone of the concept web. The learning rate was set to $0.04$ to control the speed and stability of this training process.
Crucially, the two-stage learning dynamic is not an artifact of fine-tuned parameters but a robust emergent property of the model. Repeated experiments show that this behavior—a fast-learning phase followed by a slow integration plateau, and the corresponding V-shaped response length—consistently emerges under a broad condition: when the number of Q-A pairs, $N_{QA}$, is in the regime of $1\ll N_{QA}\ll N$. This dynamic is qualitatively different from the single Q-A pair ($N_{QA}=1$) case, which produces a simple sigmoidal learning transition (the LaC phenomenon). The condition is essential: if $N_{QA}$ is too small, there is no systemic competition to drive the formation of a unified web; if it is too large, the conflicting reinforcement signals prevent any coherent structure from forming. The robust emergence of these LLM-like signatures in this intermediate regime confirms that our CoNet setup successfully isolates the core principles of structural self-organization under multi-task reinforcement.
@@ Appendix F Structural Evolution of the Concept Web in CoNet
This section provides direct, quantitative evidence from the CoNet model for the two-stage learning dynamic—a fast acquisition phase followed by a slow integration phase—that underlies our theory. Fig. 9 serves as the central exhibit for this analysis, tracking two key structural properties over the course of training: the total number of disconnected solution clusters (the “skill islands”) and the size of the single largest cluster, namely the concept web.
Figure 9: The Formation of the Concept Web in CoNet. This figure illustrates the evolution of network clusters during training, showing a structural transition from isolated skills to a unified conceptual web. In the initial phase, the Cluster Number (orange curve) spikes, indicating the rapid discovery of numerous disconnected “skill islands”. The peak marks a critical transition point. Following this peak, the Cluster Number steadily declines as these islands begin to coalesce. This merging process is confirmed by the corresponding monotonic growth in the Max Cluster Size (blue curve), which represents the formation of a single giant component. By the end of training, the network is dominated by this large, interconnected structure.
@@
The transition to the slow-learning phase occurs at the critical juncture marked by the peak of the orange curve around step 20. At this point, the system has exhausted the “low-hanging fruit” of simple, isolated solutions and has reached a state of maximal fragmentation. This peak signals the onset of the “maximally frustrated state” discussed in the main text. The subsequent, gradual decline of the “Cluster Number” curve is the direct signature of this frustrated state, representing the arduous and competitive process of integrating existing islands rather than discovering new ones. The “Max Cluster Size” (blue curve), which begins its steep, sustained ascent at precisely this moment, confirms this shift from local discovery to a global integration objective.
@@ Appendix G Learning by Phase Transition in the Slow-Learning Stage
Figure 10: Per-Problem Critical Dynamics During the Slow-Learning Stage. These plots reveal the microscopic learning dynamics for a representative subset of problems that successfully learn during the slow-learning stage, providing evidence for localized phase transitions. (Left) Accuracy trajectories for these problems. All problems shown exhibit a sharp, sigmoidal-like increase in accuracy late in the training process, characteristic of a critical learning event. (Right) Corresponding variance of the solution path length. A sharp, pronounced peak in variance directly coincides with the rapid accuracy gain for each problem. This dual signature is the hallmark of a continuous phase transition, suggesting that the slow-learning stage in CoNet is composed of discrete, sharp learning events rather than a uniform, gradual process.
@@ H.2 Expeimental Results
The algorithm consists of a strategically timed Supervised Fine-Tuning (SFT) phase, followed by a resumption of standard Reinforcement Learning from Verifiable Rewards (RLVR) training. The SFT intervention is applied at training step 50, the empirically determined point of maximum frustration where the system transitions from forming disconnected “skill islands” to integrating them. At this stage, we first identify all problems with an accuracy below 0.1, select a total of 50 low-accuracy problems, and sample a known successful reasoning path for each. This targeted adjustment serves as the “heating” step in our simulated annealing analogy; it gently increases the discoverability of latent correct paths to restore the system’s exploratory capacity. Following this brief SFT, standard RLVR training is resumed, which functions as the “cooling” phase, allowing the now more exploratory policy to settle into a new, more globally optimal configuration.
The success of this SFT-RLVR cycle is confirmed by the results in Fig. 11. The line plot on the right shows the immediate effect of the intervention: best@1 accuracy (a proxy for exploitation) temporarily dips, while best@16 accuracy (a proxy for exploration) surges. This demonstrates a successful trade of greedy exploitation for enhanced exploration, and the long-term benefit is clear, as the final annealed policy significantly surpasses the baseline RLVR policy. The histogram on the left reveals the underlying mechanism for this improvement. The SFT-RLVR cycle successfully resolves the frustration bottleneck by dramatically reducing the population of completely unsolved problems (Accuracy = 0.0) and correspondingly increasing the number of fully mastered problems (Accuracy = 1.0) compared to the standard model.
@@ I.1 Base Model and GRPO Setup
We used the DeepSeek-R1-Distill-Qwen-1.5B model (Guo et al., 2025), trained on the DeepScaleR-Preview-Dataset (Luo et al., 2025) with their modified version of the veRL open-source engine (Sheng et al., 2024; ByteDance Seed Team and verl community, ). Our Generative Reward Policy Optimization (GRPO) hyperparameters followed the first-stage settings, which specified a response length of 8192, as outlined in the DeepScaleR paper (Luo et al., 2025).
@@ I.2 SFT Dataset for Annealed-RLVR
The Supervised Fine-Tuning (SFT) dataset for the annealed-RLVR experiment (labeled “Annealed” in Fig. 6) was constructed to focus on difficult problems. First, we trained the base model using GRPO for 100 steps to create a checkpoint (CKPT100), which is in the vicinity of the maximally frustrated state. We then used this checkpoint to generate 8 rollouts for each of the approximately 40,000 problems in the training set. For the roughly 10,000 problems that yielded no correct responses, we generated an additional 52 rollouts, bringing the total to 60 per problem. From this data, we created the SFT dataset by selecting all correct responses for problems where the initial model accuracy was below 10%, which yielded 3,933 trajectories. Finally, we fine-tuned the model on these trajectories for two epochs with a learning rate of $3\times 10^{-5}$. The resulting model served as the new starting point for continuing GRPO training up to step 1,700.

### v2

@@ Appendix E CoNet Implementation Details
We configured CoNet with a specific set of structural and learning parameters. The underlying concept space was instantiated as a directed random regular graph with $N=800$ nodes, where each node has a uniform out-degree of $k=40$. In this graph, we established a multi-task learning environment consisting of $128$ randomly generated Q-A node pairs. The transition parameters $\theta_{ij}$ for all edges were initialized randomly. The use of a multi-task setting with 128 concurrent Q-A pairs is the fundamental mechanism that induces structural competition. While a single task would only require finding one efficient path, the presence of numerous tasks provides competing reinforcement signals. When multiple Q-A paths need to traverse a common node, they may “disagree” on the optimal subsequent step, leading to conflicting policy updates. This microscopic conflict is the origin of the macroscopic “frustration” discussed in the main text, compelling the system to find a shared, globally coherent structure that can arbitrate these conflicts, rather than settling for 128 disconnected “skill islands”.
@@
This advantage-based update functions as a competitive reinforcement mechanism, strengthening above-average paths and suppressing underperforming ones. This “winner-take-more” dynamic acts as an algorithmic sculpting tool, systematically pruning the vast majority of the initial $k=40$ edges from each node to carve out the sparse, efficient backbone of the concept web. The learning rate was set to $0.04$ to control the speed and stability of this training process.
Crucially, the two-stage learning dynamic is not an artifact of fine-tuned parameters but a robust emergent property of the model. Repeated experiments show that this behavior—a fast-learning phase followed by a slow integration plateau, and the corresponding V-shaped response length—consistently emerges under a broad condition: when the number of Q-A pairs, $N_{QA}$, is in the regime of $1\ll N_{QA}\ll N$. This dynamic is qualitatively different from the single Q-A pair ($N_{QA}=1$) case, which produces a simple sigmoidal learning transition (the LaC phenomenon). The condition is essential: if $N_{QA}$ is too small, there is no systemic competition to drive the formation of a unified web; if it is too large, the conflicting reinforcement signals prevent any coherent structure from forming. The robust emergence of these LLM-like signatures in this intermediate regime confirms that our CoNet setup successfully reproduces the essential dynamics of structural self-organization under multi-task reinforcement.
@@ Appendix F Structural Evolution of the Concept Web in CoNet
This section provides direct, quantitative evidence from the CoNet model for the two-stage learning dynamic—a fast acquisition process followed by a slow integration phase—that underlies our theory. Fig. 9 serves as the central exhibit for this analysis, tracking two key structural properties over the course of training: the total number of disconnected solution clusters (the “skill islands”) and the size of the single largest cluster, namely the concept web.
Figure 9: The Formation of the Concept Web in CoNet. This figure illustrates the evolution of network clusters during training, showing a structural reorganization from isolated skills to a unified conceptual web. In the initial phase, the Cluster Number (orange curve) spikes, indicating the rapid discovery of numerous disconnected “skill islands”. The peak marks a critical structural juncture. Following this peak, the Cluster Number steadily declines as these islands begin to coalesce. This merging process is confirmed by the corresponding monotonic growth in the Max Cluster Size (blue curve), which represents the formation of a single giant component. By the end of training, the network is dominated by this large, interconnected structure.
@@
The shift to the slow-learning phase occurs at the critical juncture marked by the peak of the orange curve around step 20. At this point, the system has exhausted the “low-hanging fruit” of simple, isolated solutions and has reached a state of maximal fragmentation. This peak signals the onset of the “maximally frustrated state” discussed in the main text. The subsequent, gradual decline of the “Cluster Number” curve is the direct signature of this frustrated state, representing the arduous and competitive process of integrating existing islands rather than discovering new ones. The “Max Cluster Size” (blue curve), which begins its steep, sustained ascent at precisely this moment, confirms this shift from local discovery to a global integration objective.
@@ Appendix G Learning by Phase Transition in the Slow-Learning Stage
Figure 10: Per-Problem Critical Dynamics During the Slow-Learning Stage. These plots reveal the microscopic learning dynamics for a representative subset of problems that successfully learn during the slow-learning stage, providing evidence for localized phase transitions. (Left) Accuracy trajectories for these problems. All problems shown exhibit a sharp, sigmoidal-like increase in accuracy rate in the training process, characteristic of a critical learning event. (Right) Corresponding variance of the solution path length. A sharp, pronounced peak in variance directly coincides with the rapid accuracy gain for each problem. This dual signature is the hallmark of a continuous phase transition, suggesting that the slow-learning stage in CoNet is composed of discrete, sharp learning events rather than a uniform, gradual process.
@@ H.2 Expeimental Results
Figure 11: Structural Impact of Annealed-RLVR on the Concept Web. Evolution of concept web for standard RLVR (blue) versus Annealed-RLVR (orange). The SFT intervention at step 50 (dashed line) induces an immediate drop in the Annealed model’s cluster size. Subsequently, the Annealed model recovers and surpasses the standard RLVR baseline, which exhibits slower growth, ultimately forming a larger final concept web.
@@
The algorithm consists of a strategically timed SFT phase, followed by a resumption of standard RLVR training. The SFT intervention is applied at training step 50, the empirically determined state of maximum frustration, marking the crossover from the formation of disconnected “skill islands” to their subsequent integration. At this stage, we first identify all problems with an accuracy below 0.1, select a total of 50 low-accuracy problems, and sample a known successful reasoning path for each. This targeted adjustment serves as the “heating” step in our simulated annealing analogy; it gently increases the discoverability of latent correct paths to restore the system’s exploratory capacity. Following this brief SFT, standard RLVR training is resumed, which functions as the “cooling” phase, allowing the now more exploratory policy to settle into a new, more globally optimal configuration.
Fig. 11 provides the direct structural evidence for this annealing process. As shown by the orange “Annealed” curve, the SFT intervention at step 50 locally disrupted network connections, resulting in a decline in the concept web size. However, this disruption is precisely what allows the system to overcome the integration bottleneck. While the standard RLVR model (blue curve) is shown to grow slowly, trapped in the maximally frustrated state, the Annealed model’s cluster size rapidly recovers and grows to a larger final scale. This demonstrates that the SFT intervention, by resolving the competitive frustration, enables the subsequent RLVR “cooling” phase to successfully expand the network and integrate a greater number of “skill islands”.
The success of this SFT-RLVR cycle is also confirmed by the results in Fig. 12. The line plot on the right shows the immediate effect of the intervention: best@1 accuracy (a proxy for exploitation) temporarily dips, while best@16 accuracy (a proxy for exploration) surges. This demonstrates a successful trade of greedy exploitation for enhanced exploration, and the long-term benefit is clear, as the final annealed policy significantly surpasses the baseline RLVR policy. The histogram on the left reveals the underlying mechanism for this improvement. The SFT-RLVR cycle successfully resolves the frustration bottleneck by reducing the population of completely unsolved problems (Accuracy = 0.0) and correspondingly increasing the number of fully mastered problems (Accuracy = 1.0) compared to the standard model.
@@ I.1 Base Model and GRPO Setup
We used the DeepSeek-R1-Distill-Qwen-1.5B model (Guo et al., 2025), trained on the DeepScaleR-Preview-Dataset (Luo et al., 2025a) with their modified version of the veRL open-source engine (Sheng et al., 2024; ByteDance Seed Team and verl community, ). Our Group Relative Policy Optimization (GRPO) hyperparameters followed the first-stage settings, which specified a response length of 8192, as outlined in the DeepScaleR paper (Luo et al., 2025a).
@@ I.2 SFT Dataset for Annealed-RLVR
The SFT dataset for the annealed-RLVR experiment (labeled “Annealed” in Fig. 6) was constructed to focus on difficult problems. First, we trained the base model using GRPO for 100 steps to create a checkpoint (CKPT100), which is in the vicinity of the maximally frustrated state. We then used this checkpoint to generate 8 rollouts for each of the approximately 40,000 problems in the training set. For the roughly 10,000 problems that yielded no correct responses, we generated an additional 52 rollouts, bringing the total to 60 per problem. From this data, we created the SFT dataset by selecting all correct responses for problems where the initial model accuracy was below 10%, which yielded 3,933 trajectories. Finally, we fine-tuned the model on these trajectories for two epochs with a learning rate of $3\times 10^{-5}$. The resulting model served as the new starting point for continuing GRPO training up to step 1,700.
@@ I.3 SFT Dataset for Catastrophic Forgetting
I.4 Computational Cost Analysis
The overall computational complexity of Annealed-RLVR is dominated by the iterative GRPO training, making it nearly identical to the standard RLVR baseline. The only addition is a one-time cost for the SFT intervention, which is applied once at the point of maximal frustration (around step 100 in our case). This one-time cost consists of two phases:
  1. 1.
    Evaluation Pass: An evaluation pass over the $N$ problems in the training set (where $N\approx 40,000$) is required to identify the low-accuracy problems. As described in Section I.2, this involved generating 8 to 60 rollouts per problem. This cost can be formally described as $O(N\cdot k)$ for a fixed $k$ rollouts, or $O(N\cdot\log(k))$ if using an adaptive, binary-like rollout strategy to find problems below the accuracy threshold.
  2. 2.
    SFT Phase: The SFT itself is performed on the small subset of resulting correct trajectories. Letting $\eta$ be the accuracy threshold parameter (e.g., $\eta=0.1$ for our 10% threshold), the SFT cost is proportional to $O(\eta N)$. As detailed in Section I.2, this amounted to only 3,933 trajectories (approximately $\eta N$) in our experiment.
This $O(\eta N)$ SFT cost is significantly smaller than the evaluation cost. Both of these marginal, one-time costs are much smaller than the total iterative cost of the full GRPO training run (e.g., 1,700 steps).
@@ I.4 Computational Cost Analysis
(a) Training Sample (512 Problems)
(b) Minerva
(c) AIME 2024/2025
Figure 13: Mechanism of Annealed-RLVR: Accuracy Distributions. This figure complements the best@k curves in the main text by revealing the underlying mechanism of performance improvement. It shows the per-problem accuracy histograms for the checkpoints of Annealed-RLVR ("Annealed") versus the standard RLVR ("RLVR") baseline, evaluated on (a) the in-distribution set (Random 512), (b) the OOD Minerva dataset, and (c) the OOD AIME 2024/2025 datasets. The histograms demonstrate that the Annealed-RLVR intervention systematically reduces the population of unsolved problems (low-accuracy bins) and increases the population of mastered problems (high-accuracy bins).
@@
Dataset: AIME 2024 & 2025
best@1 GRPO 20.11(0.10) 23.08(0.11) 24.04(0.12) 26.98(0.12) 27.77(0.11) 28.11(0.11) 28.10(0.11)
Annealed 23.27(0.11 23.86(0.12 26.77(0.12 27.67(0.11 28.26(0.11 27.93(0.11
best@4 GRPO 33.31(0.17) 38.44(0.19) 40.33(0.19) 42.74(0.18) 43.36(0.19) 43.51(0.18) 43.88(0.18)
Annealed 39.36(0.19) 40.21(0.19) 43.38(0.19) 43.76(0.19) 43.79(0.19) 43.42(0.19)
best@16 GRPO 45.3(0.3) 52.4(0.3) 54.4(0.3) 55.5(0.3) 56.9(0.3) 56.1(0.3) 57.1(0.3)
Annealed 54.1(0.3) 54.7(0.3) 57.6(0.3) 57.7(0.3) 58.1(0.3) 57.8(0.3)
best@64 GRPO 55.9(0.5) 62.4(0.5) 64.2(0.5) 64.5(0.6) 65.9(0.6) 64.8(0.6) 66.2(0.6)
Annealed 63.8(0.5) 64.2(0.5) 67.0(0.5) 66.7(0.5) 67.5(0.5) 67.7(0.5)